# Catching the Details: Self-Distilled RoI Predictors for Fine-Grained MLLM Perception

### v2

@@ 4 Experiments
Table 2: Performance on Vision-Centric and High-Resolution benchmarks. $\dagger$ denotes results reported in the original publication..
@@
Benchmark Settings. To ensure a comprehensive evaluation, we test our SD-RPN framework across a range of benchmarks, following the protocols of established works (Shi et al., 2025; Zhang et al., 2025; Zheng et al., 2025). Our results are presented in two parts to clearly delineate performance across different domains. The first part focuses on the Document & OCR category, which assesses performance on text-rich images, and includes five benchmarks: DocVQA (Mathew et al., 2020), ChartQA (Masry et al., 2022), OCRBench (Liu et al., 2024b), InfoVQA (Mathew et al., 2022), and TextVQA (Singh et al., 2019). The second part focuses on Vision-Centric and High-Resolution tasks, including V-Star Bench (V*) (Wu & Xie, 2023), POPE (Li et al., 2023), and HR-Bench (Wang et al., 2024b). Notably, for TextVQA, we do not provide the model with external OCR annotations, a strategy adopted from ViCrop (Zhang et al., 2025) to ensure a fair evaluation of the MLLM’s intrinsic perceptual capabilities.
@@
Baselines and Comparison Methods. To evaluate SD-RPN, we integrate it into three prominent MLLM families. Our primary experiments are conducted on the widely-used LLaVA-1.5 (Liu et al., 2023a) architecture, across both its 7B and 13B scales. To demonstrate the generalizability of our approach, we also apply it to the more recent DeepSeek-VL (Lu et al., 2024) and Qwen2.5-VL (Bai et al., 2025) models. We compare our method against several representative techniques for high-resolution visual processing. For the LLaVA-1.5 models, we benchmark against S2 (Shi et al., 2024), a full-tuning method, and ViCrop (Zhang et al., 2025), a training-free cropping baseline. For the Qwen2.5-VL model, we additionally compare against DeepEyes (Zheng et al., 2025), a recent reinforcement learning-based method. To ensure a rigorous and fair comparison for full reproducibility, we have re-implemented all competing methods within the unified lmms-eval evaluation library (Zhang et al., 2024a), unless noted otherwise.
@@
Implementation Details. To train our proposed SD-RPN, we generate pseudo-labels from a combined dataset of GQA (Hudson & Manning, 2019) and OCR-VQA (Mishra et al., 2019), which provides supervision for natural scenes and text-rich images, respectively. These pseudo-labels are derived from the models’ internal response-to-image attention maps. Inspired by recent analyses of MLLM attention mechanisms (Kang et al., 2025b), we select attention maps from the middle layers for this process. During the pseudo-label generation pipeline, the relative attention thresholds for defining the foreground ($\mathcal{S}_{fg}$) and background ($\mathcal{S}_{bg}$) sets are empirically set to 0.2 and 0.1, respectively. For experiments involving Qwen2.5-VL, which natively supports dynamic resolutions, we standardize the input for consistency. On most benchmarks, we set the number of visual tokens to 576 to align with the LLaVA and DeepSeek-VL baselines. For high-resolution benchmarks, we increase this to a maximum of 4096 visual tokens to effectively process the detailed inputs. Further details on the layer selection rationale and hyper-parameters can be found in Appendix A.
@@ 4.2 Main Results
We present our main results in Tab. 1 for text-rich benchmarks and Tab. 2 for vision-centric benchmarks. Our evaluation assesses both performance and efficiency. In addition to these quantitative results, we provide a qualitative comparison in Appendix D.
Performance. Integrating our SD-RPN framework yields substantial and consistent performance gains across all evaluated MLLM families. On the text-rich benchmarks (Tab. 1), our method boosts the average score by approximately 7% over the LLaVA-1.5 and DeepSeek-VL baselines and 3% over the more powerful Qwen2.5-VL baseline. The benefits are also evident on vision-centric and high-resolution tasks (Tab. 2), where our approach achieves average improvements of over 10% on V* Bench and over 6% on HR-Bench. Notably, on high-resolution benchmarks, SD-RPN achieves performance comparable or superior to DeepEyes which relies visual Chain-of-Thought (CoT) that incurs substantially greater computational cost. To ensure a fair comparison on V-Star Bench, we note that the image cropping strategy employed by ViCrop was disabled during our evaluation.
@@
4.3 Analysis of the Performance-Throughput Trade-off
@@ 5 Conclusion
We proposed SD-RPN, a self-distilled region proposal framework that efficiently exploits the intrinsic localization signals of MLLMs to identify Regions of Interest without external annotations or auto-regressive decoding. By attaching a lightweight RPN to frozen backbones and training it with denoised pseudo-labels, our method achieves consistent improvements in fine-grained perception while maintaining strong efficiency and generalization. Extensive experiments confirm its advantage over both full-image scaling and training-free heuristics, and ablations highlight its robustness and data efficiency. Our work establishes a principled direction for scalable high-resolution perception in MLLMs, and opens avenues toward adaptive token allocation and broader multimodal applications such as video and document understanding.
@@
Algorithm Details. To provide a concrete and reproducible specification of our training procedure, we detail the complete end-to-end process in Algorithm 1. The algorithm consists of a main training loop and the core helper routines. The main training loop iterates through the dataset, performing two key stages for each sample. First, the ”teacher” stage (lines 3-10) uses the full, frozen MLLM to generate a pseudo-label. This involves auto-regressively generating a response, extracting the raw response-to-image attention map from a pre-selected middle layer, and then refining this map using our RemoveSinkTokens and AssignLabels helper functions. This produces the final denoised pseudo-label, $\bar{\mathbf{M}}_{\text{RoI}}$. Second, the ”student” stage (lines 11-15) trains the lightweight RPN. The RPN performs a forward pass on the original input to predict a dense RoI map $\hat{\mathbf{M}}_{\text{RoI}}$. A masked BCEWithLogits loss is then computed between the predicted map and the pseudo-label target, ensuring that ambiguous regions (where the mask is false) do not contribute to the gradient. Finally, the RPN’s parameters are updated via backpropagation.
@@ Appendix B More Ablations
Appendix C Prompt Usage
@@ Appendix D Visualization
Appendix E Why Predicting RoI Scores Outperforms Using Raw Attention
@@
A lightweight RoI head trained to regress $A$ learns $\mathbb{E}[A\mid X]$, which (i) is order-preserving with $\eta(x)$ by the affine forms above, and (ii) enjoys reduced variance. Thresholding a learned predictor $\hat{h}(X)$ therefore yields more accurate and stable RoI proposals than thresholding raw attention $A$, matching our empirical gains. Our proposed Label Assignment strategy can be seen as a further refinement of this principle, where we structure the learning problem as a classification task on high-confidence tokens to make the training process even more robust to the noise in $A$.
Appendix F The Use of Large Language Models

### v3

@@ 4 Experiments
Table 2: Performance on Vision-Centric and High-Resolution benchmarks. $\dagger$ denotes results reported in the original publication.
@@
Benchmark Settings. To ensure a comprehensive evaluation, we test our SD-RPN framework across a range of benchmarks, following the protocols of established works (Shi et al., 2025a; Zhang et al., 2025a; Zheng et al., 2025). Our results are presented in two parts to clearly delineate performance across different domains. The first part focuses on the Document & OCR category, which assesses performance on text-rich images, and includes five benchmarks: DocVQA (Mathew et al., 2020), ChartQA (Masry et al., 2022), OCRBench (Liu et al., 2024b), InfoVQA (Mathew et al., 2022), and TextVQA (Singh et al., 2019). The second part focuses on Vision-Centric and High-Resolution tasks, including V-Star Bench (V*) (Wu and Xie, 2023), POPE (Li et al., 2023), and HR-Bench (Wang et al., 2024b). Notably, for TextVQA, we do not provide the model with external OCR annotations, a strategy adopted from ViCrop (Zhang et al., 2025a) to ensure a fair evaluation of the MLLM’s intrinsic perceptual capabilities.
@@
Baselines and Comparison Methods. To evaluate SD-RPN, we integrate it into three prominent MLLM families. Our primary experiments are conducted on the widely-used LLaVA-1.5 (Liu et al., 2023a) architecture, across both its 7B and 13B scales. To demonstrate the generalizability of our approach, we also apply it to the more recent DeepSeek-VL (Lu et al., 2024) and Qwen2.5-VL (Bai et al., 2025) models. We compare our method against several representative techniques for high-resolution visual processing. For the LLaVA-1.5 models, we benchmark against S2 (Shi et al., 2024), a full-tuning method, and ViCrop (Zhang et al., 2025a), a training-free cropping baseline. For the Qwen2.5-VL model, we additionally compare against DeepEyes (Zheng et al., 2025) and Thyme (Zhang et al., 2025b), which are recent reinforcement learning-based method. To ensure a rigorous and fair comparison for full reproducibility, we have re-implemented all competing methods within the unified lmms-eval evaluation library (Zhang et al., 2024a), unless noted otherwise.
@@
Implementation Details. To train our proposed SD-RPN, we generate pseudo-labels from a combined dataset of GQA (Hudson and Manning, 2019) and OCR-VQA (Mishra et al., 2019), which provides supervision for natural scenes and text-rich images, respectively. These pseudo-labels are derived from the models’ internal response-to-image attention maps. Inspired by recent analyses of MLLM attention mechanisms (Kang et al., 2025b), we select attention maps from the middle layers for this process. During the pseudo-label generation pipeline, the relative attention thresholds for defining the foreground ($\mathcal{S}_{fg}$) and background ($\mathcal{S}_{bg}$) sets are empirically set to 0.2 and 0.1, respectively. For experiments involving Qwen2.5-VL, which natively supports dynamic resolutions, we standardize the input for consistency. On most benchmarks, we set the number of visual tokens to 576 to align with the LLaVA and DeepSeek-VL baselines. For high-resolution benchmarks, we increase this to a maximum of 4096 visual tokens to effectively process the detailed inputs. Further details on the layer selection rationale and hyper-parameters can be found in Appendix A.
@@ 4.2 Main Results
We present our main results in Tab. 1 for text-rich benchmarks and Tab. 2 for vision-centric benchmarks. Our evaluation assesses both performance and efficiency.
Performance. Integrating our SD-RPN framework yields substantial and consistent performance gains across all evaluated MLLM families. On the text-rich benchmarks (Tab. 1), our method boosts the average score by approximately 7% over the LLaVA-1.5 and DeepSeek-VL baselines and 3% over the more powerful Qwen2.5-VL baseline. The benefits are also evident on vision-centriand high-resolution tasks (Tab. 2), where our approach achieves average improvements of over 10% on V* Bench and over 6% on HR-Bench. Notably, on high-resolution benchmarks, SD-RPN achieves performance comparable or superior to DeepEyes which relies visual Chain-of-Thought (CoT) that incurs substantially greater computational cost. To ensure a fair comparison on V-Star Bench, we note that the image cropping strategy employed by ViCrop was disabled during our evaluation. Beyond these quantitative results, Fig. 5 offers a qualitative comparison with the LLaVA-1.5 and Qwen2.5-VL baselines. Additional visual examples are provided in Appendix E.
@@
Figure 5: Qualitative comparisons on TextVQA and V-Star benchmarks. These examples show challenging cases where the baseline model fails while our SD-RPN succeeds. For each question, our model generates a dense RoI Heatmap (center) to localize the relevant area. The Mixed Image (right) shows the final high-confidence region identified by SD-RPN, which is used for fine-grained perception to find the correct answer.
@@ 4.3 Ablation Study
In this section, we first present ablations for our core design choices, including the pseudo-label generation process, backbone layer depth, and data efficiency. These experiments were conducted using default thresholds ($\tau_{fg}=0.2,\tau_{bg}=0.1$). We then provide a detailed ablation study on these $\tau_{fg}$ and $\tau_{bg}$ thresholds to determine their optimal values. [^1] We thank the reviewers for motivating this deeper analysis of our hyperparameter settings.
@@
Table 5: Ablation study on pseudo-label thresholds $\tau_{fg}$ and $\tau_{bg}$.
$\bm{\tau_{fg}}$ $\bm{\tau_{bg}}$
OCR-
Bench
Text-
VQA
POPE V* Ave.
Baseline 31.4 46.1 85.9 50.3 53.4
0.10 0.10 35.1 57.4 87.4 65.4 61.3
0.15 0.10 36.3 57.9 87.2 67.5 62.2
0.20 0.10 37.0 58.7 87.1 67.5 62.6
0.25 0.10 37.0 58.6 86.7 68.1 62.6
0.30 0.10 36.6 58.1 86.9 64.9 61.6
0.20 0.075 36.7 58.4 86.9 70.7 63.2
0.20 0.05 37.3 58.8 87.2 70.7 63.5
0.20 0.03 37.8 58.7 87.6 68.6 63.2
@@
On Foreground and Background Threshold. Tab. 5 ablates the pseudo-label thresholds. We find that setting $\tau_{fg}=\tau_{bg}=0.10$ (a simple binary classification) performs poorly (61.3% avg.), as it suffers from ambiguous tokens. By optimizing the thresholds to $\tau_{fg}=0.20$ and $\tau_{bg}=0.05$, we establish a new, stronger default. This setting achieves a 63.5% average performance, surpassing our original default (62.6%) and the baseline (53.4%) by 0.9% and 10.1%, respectively.
@@ 5 Conclusion
We proposed SD-RPN, a self-distilled region proposal framework that efficiently exploits the intrinsic localization signals of MLLMs to identify Regions of Interest without external annotations or auto-regressive decoding. By attaching a RPN to frozen backbones and training it with denoised pseudo-labels, our method achieves consistent improvements in fine-grained perception while maintaining strong efficiency and generalization. Extensive experiments confirm its advantage over both full-image scaling and training-free heuristics, and ablations highlight its robustness and data efficiency. Our work establishes a principled direction for scalable high-resolution perception in MLLMs, and opens avenues toward adaptive token allocation and broader multimodal applications such as video and document understanding.
6 Acknowledgements
This work was supported in part by the Start-up Grant (No. 9610680) of the City University of Hong Kong, Young Scientist Fund (No. 62406265) of NSFC, and the Australian Research Council under Projects DP240101848 and FT230100549.
@@ 6 Acknowledgements
Algorithm Details. To provide a concrete and reproducible specification of our training procedure, we detail the complete end-to-end process in Algorithm 1. The algorithm consists of a main training loop and the core helper routines. The main training loop iterates through the dataset, performing two key stages for each sample. First, the ”teacher” stage (lines 3-10) uses the full, frozen MLLM to generate a pseudo-label. This involves auto-regressively generating a response, extracting the raw response-to-image attention map from a pre-selected middle layer, and then refining this map using our RemoveSinkTokens and AssignLabels helper functions. This produces the final denoised pseudo-label, $\bar{\mathbf{M}}_{\text{RoI}}$. Second, the ”student” stage (lines 11-15) trains the RPN. The RPN performs a forward pass on the original input to predict a dense RoI map $\hat{\mathbf{M}}_{\text{RoI}}$. A masked BCEWithLogits loss is then computed between the predicted map and the pseudo-label target, ensuring that ambiguous regions (where the mask is false) do not contribute to the gradient. Finally, the RPN’s parameters are updated via backpropagation.
@@ Appendix B More Ablations
These experiments blow were conducted using default thresholds ($\tau_{fg}=0.2,\tau_{bg}=0.1$), as they were implemented before threshold ablation.
@@
Appendix C More Comparison and Analysis
C.1 Analysis of the Performance-Throughput Trade-off
C.2 Comparison to External Grounding Models
A practical concern regarding our SD-RPN is its computational cost, as it utilizes a portion of the MLLM. It may seem counter-intuitive, but our SD-RPN is, in fact, significantly more computationally efficient than using an external grounding model like Grounding-DINO or YOLO-World. The reason is that our RPN re-uses computations that the MLLM must perform anyway. An external model (e.g., Grounding-DINO) would require redundant computation in the form of a separate, full image encoding pass, which is computationally expensive. In contrast, our approach uses feature re-use: our RPN attaches to the MLLM’s middle layers and operates on features that are already computed during the MLLM’s initial forward pass. The only extra cost is passing these features through the RPN layers (e.g., 3 layers in our B15R3 configuration), which is a minimal addition.
To provide a concrete analysis, we benchmarked our method (on LLaVA-1.5-7B) against a Grounding-DINO-Base model on a NVIDIA A6000 GPU. The results, presented in Table 10, are clear. Our RPN is over 20x faster in its specific prediction step (10.3ms vs 231.2ms) because it avoids a redundant, full-image encoding. This also leads to a much faster overall evaluation time on the V* benchmark. Our method is also significantly more accurate (62.6% vs 57.6% avg.). This is because our RPN is an integrated part of the MLLM and has access to its contextual understanding of the query. External models are excellent at finding literal objects (e.g., ”the bottom man”) but fail at abstract, reasoning-based queries that don’t name a specific object, such as the example in the last row of Figure 7. Our RPN, trained on the MLLM’s internal attention, excels at this.
Table 10: Efficiency and Performance: Ours (SD-RPN) vs. External Grounding-DINO. The ”Grounding-DINO-B (1Box)” method crops one sub-image using the single most confident bounding box, while ”(2Box)” crops two sub-images using the top 2 boxes.
Method RPN Time (ms) V* Eval Time (s) OCRBench TextVQA POPE V* Ave.
Ours (SD-RPN) 10.3 73 37.0 58.7 87.1 67.5 62.6
Grounding-DINO-B (2Box) 231.2 248 32.5 54.5 82.4 62.3 57.6
Grounding-DINO-B (1Box) 231.2 113 32.5 52.6 86.9 56.5 57.1
C.3 Deeper Analysis on Sources of Attention Noise
In our work, we identify and actively mitigate two primary sources of noise from the raw attention maps used to generate pseudo-labels. First, there are Type 1: Sink Tokens. We build on recent studies (Darcet et al., 2023; Kang et al., 2025a) which identify ”sink tokens”—visual tokens that attract high attention despite being semantically irrelevant to the grounded object. Second, there are Type 2: Ambiguous Foreground/Background Signals. This is a more common, semantic, and task-dependent form of noise. As illustrated in Fig. 3 and described in Section 3.2, these are ”noisy maps” that exhibit ”erroneously high attention in background areas and incomplete activation across the true foreground object.”
As the first type of sink token belongs to a well-defined source of noise, we now investigate the second type. Our empirical study finds that questions or images in the following three categories are prone to producing noisy attention maps, as illustrated in Fig. 7. The first source of noise is cluttered scenes or dense text (e.g., the first sample in Fig. 7); in these cases, the model’s attention can be fraught with noise as it diffuses across multiple irrelevant-but-plausible regions. The second type of noise is caused by abstract or reasoning-heavy queries. In practice, we find that a simple query (e.g., ”find the cat”) tends to produce clean attention. However, many VQA queries are abstract, such as the second sample in Fig. 7. For such queries, the attention map reflects the model’s complex reasoning process, which can appear noisy as it attends to multiple pieces of related (but not all relevant) information before settling on an answer. The last type of noise is related to incomplete activation, where the attention map only weakly or partially covers the true region of interest, as shown in the last sample of Fig. 7.
Figure 7: Examples of noisy and sparse attention maps for abstract queries.
C.4 Comparison with Token Compression Methods
We provide a detailed comparison of our SD-RPN with other SoTA methods designed for high-resolution visual understanding in this section. To facilitate a fair and transparent comparison, Table 11 summarizes the performance of our method against several competitive models on key document and OCR benchmarks. The table also highlights critical architectural and data differences, such as the base Large Language Model (LLM), the size of the training dataset, and the maximum input resolution, all of which are significant factors that influence final performance.
Table 11: Comparison with SoTA high-resolution methods on document and OCR benchmarks. #Denotes results obtained using OCR tokens. *Denotes results reported in the LLaVA-TokenPacker-HD paper (Li et al., 2025). The ”+152K” in the #Data column indicates the number of additional samples used for fine-tuning our SD-RPN on the corresponding baseline models.
Method LLM #Data Max Res. TextVQAval OCRBenchtest DocVQAtest ChartQAtest
LLaVA-1.5 Vicuna-7B 1.2M $\times336$ 58.2# 31.4 22.2 18.1
LLaVA-1.5 + SD-RPN Vicuna-7B +152K $\times336$ 64.2# 37.0 35.1 20.1
Monkey (Li et al., 2024b) Qwen-7B 1.4M $\times896$ 67.6 - 66.5 65.1
LLaVA-NeXT (Liu et al., 2024a) Vicuna-7B 1.3M $\times672$ 64.9* - - -
[…]
LLaVA-TokenPacker-HD (Li et al., 2025) Vicuna-7B 2.7M $\times1088$ 68.0* 452* 60.2* -
DeepSeek-VL Vicuna-7B - $\times1024$ 63.0 42.4 49.3 61.4
DeepSeek-VL + SD-RPN Vicuna-7B +152K $\times1024$ 71.5 47.9 64.5 63.6
Qwen2.5-VL Qwen2.5-7B - $\times672$ 81.1 81.5 92.6 82.6
Qwen2.5-VL + SD-RPN Qwen2.5-7B +152K $\times672$ 83.5 82.9 94.4 85.5
The results in Table 11 lead to two primary conclusions. First, the overall performance of our method, like others, is strongly correlated with the capabilities of the underlying base MLLM. When applied to an earlier baseline such as LLaVA-1.5, our method’s absolute scores on benchmarks like DocVQA are understandably lower than SoTA methods built on more advanced foundations, which often leverage richer training sets and higher input resolutions. This performance difference is primarily a reflection of the base model’s capacity rather than a limitation of our enhancement module. Second, our SD-RPN module demonstrates consistent efficacy across base models of varying strengths. When integrated with a more competitive baseline like DeepSeek-VL-7B, our method achieves performance on par with or slightly exceeding other leading methods. To further validate this generalizability, we applied SD-RPN to the powerful Qwen2.5-VL model. Even on this SoTA baseline, our method delivers significant and consistent gains across all benchmarks, utilizing an identical and minimal fine-tuning recipe.
This highlights a key advantage of our approach: SD-RPN provides a data-efficient and scalable pathway for enhancing the fine-grained perceptual capabilities of MLLMs. Unlike methods that necessitate extensive supervised fine-tuning on massive datasets, our self-distillation technique offers a resource-efficient solution for improving high-resolution understanding that can readily adapt to the continuous advancements in base MLLMs.
Appendix D Prompt Usage
@@ Appendix E Visualization
Figure 9: Failure case analysis of SD-RPN on the LLaVA-1.5-7B baseline. We present examples where the baseline model succeeds, but our method fails. These failures can be categorized into three main types: incomplete activation (top-left), localization error (top-right), and spatial relationship error (bottom row).
In addition to the successful cases, we also provide a failure case analysis in Fig. 9. The failure cases can be broadly attributed to three main reasons:
  • Incomplete Activation. The top-left ”cup” case is a representative example. The RPN correctly identifies the measuring cup but only activates on a small portion of it. This partial RoI provides incomplete visual information to the MLLM, leading to an erroneous prediction.
  • Localization Error. The top-right ”book” example illustrates this type. The RPN is supposed to highlight the book’s cover to read the name but erroneously highlights an unrelated region. Given this misleading visual cue, the MLLM fails to produce the correct result.
  • Spatial Relationship Error. The bottom row shows errors of this type. Here, the RPN highlights all relevant objects correctly (e.g., ”bowl” and ”faucet”). However, the MLLM still produces the wrong result, suggesting that it struggles to preserve and reason about the precise spatial relationships between objects when they are processed as fine-grained crops.
Appendix F Why Predicting RoI Scores Outperforms Using Raw Attention
@@
A RoI head trained to regress $A$ learns $\mathbb{E}[A\mid X]$, which (i) is order-preserving with $\eta(x)$ by the affine forms above, and (ii) enjoys reduced variance. Thresholding a learned predictor $\hat{h}(X)$ therefore yields more accurate and stable RoI proposals than thresholding raw attention $A$, matching our empirical gains. Our proposed Label Assignment strategy can be seen as a further refinement of this principle, where we structure the learning problem as a classification task on high-confidence tokens to make the training process even more robust to the noise in $A$.
Appendix G The Use of Large Language Models